### v1

@@ Abstract
Precession of a converged beam during acquisition of a 4D-STEM dataset improves strain, orientation, and phase mapping accuracy by averaging over continuous angles of illumination. Precession experiments usually rely on integrated systems, where automatic alignments lead to fast, high-quality results. The dependence of these experiments on specific hardware and software is evident even when switching to non-integrated detectors on a precession tool, as experimental set-up becomes challenging and time-consuming. Here, we introduce multi-angle precession electron diffraction (MAPED): a method to perform precession electron diffraction by collecting sequential 4D-STEM scans at different incident beam tilts. The multiple diffraction datasets are averaged together post-acquisition, resulting in a single dataset that minimizes the impact of the curvature and orientation of the Ewald sphere relative to the crystal under study. Our results demonstrate that even four additional tilts improved measurement of material properties, namely strain and orientation, as compared to single-tilt 4D-STEM experiments. We show the versatility and flexibility of our few-tilt precession approach with data collected on a number of microscopes with different hardware configurations and a variety of detectors.
@@ 1 Introduction
Modern STEM experiments often include 4D-STEM, where an electron beam is rastered across a specimen, and diffraction data are collected at each position in real space creating a 4-dimensional dataset (Ophus, 2019). The 4D-STEM data can be analyzed to map properties of materials that would be inaccessible or challenging to probe with conventional methods. For 4D-STEM strain experiments, data are collected with a probe formed by a small convergence angle, which reduces overlap of diffracted beams in reciprocal space. This approach can be extended to amorphous materials, which is not possible with conventional high resolution techniques (Gammer et al., 2018; Kennedy et al., 2025). Similarly, the orientation and phase of a sample can be studied using 4D-STEM, and when coupled with computational diffraction pattern matching, enables automated crystal orientation mapping (ACOM) (Cautaerts et al., 2022; Ophus et al., 2022). Despite the immense benefits of strain mapping and ACOM techniques for studying the structure and properties of material systems, there can be practical challenges for thicker and more complex structures, due to multiple scattering of the electron beam leading to dynamical effects in the diffraction patterns (Müller et al., 2012).
@@
One of the most influential innovations for mitigating dynamical diffraction effects in S/TEM was introduced by Vincent and Midgley in 1994, who demonstrated precession electron diffraction with a convergent beam (Vincent and Midgley, 1994). This approach, which builds on earlier “hollow cone” illumination strategies  (Krakow and Howland, 1976; Kondo et al., 1984), uses rocking of the Ewald sphere to sample a wider range of diffraction space. The resulting diffraction patterns exhibit more uniform Bragg disk intensities, are less sensitive to sample mistilt, and can more easily excite higher-order reflections, making them more straightforward to index and more robust against dynamical scattering (Rouviere et al., 2013; Rauch et al., 2010). By averaging over multiple directions of incident electron beams, diffraction contrast can be suppressed, and it has even enabled the measurement of polarization through differential phase contrast (DPC) measurements (Kohno et al., 2022). Similar to this work, recent studies have implemented multiple tilts for analysis of electric polarization as an alternative to DPC imaging (Toyama et al., 2024; Flathmann et al., 2025).
@@
Precession electron diffraction can now be directly integrated into 4D-STEM workflows, where synchronized beam precession, raster scanning, and camera acquisition are handled by vendor-supplied systems. These integrated tools maintain precise alignment during acquisition, producing sharp diffraction patterns, which can be leveraged for high-quality analysis including strain mapping, orientation indexing, and digital dark field analysis (Corrêa et al., 2024; Diebold et al., 2025; MacLaren et al., 2025). These precession tools can be challenging to align, especially when they are used with hardware such as detectors and apertures that are not fully integrated in the system. Although there is an increasing emphasis and availability on precession tools, few microscopes have integrated precession equipment, making the implementation of precession techniques limited.
@@
Here we present multi-angle precession electron diffraction (MAPED): a data-acquisition and post-processing strategy based on sequentially acquired 4D-STEM datasets at five distinct beam tilts to perform precession experiments. This sequential tilting approach allows us to replicate the averaging effect of hardware-integrated precession with the benefit of flexibility across different microscopes and detectors. We found that even four additional tilts ($\sim$0.5-1) dramatically improved the accuracy of the resulting strain maps, while single-tilt datasets sometimes yielded non-physical results due to dynamical diffraction. For challenging cases, such as phase discrimination in structurally similar crystals, we also simulated silicon samples along the [110] zone at various thicknesses to determine how well disk intensities matched those predicted by structure factor calculations. To test robustness, we examined data acquired on both corrected and uncorrected TEMs, using multiple camera systems, and assessed the effect of tilt count on indexing fidelity and strain reconstruction accuracy. This work provides a generalizable and practical framework for enhancing 4D-STEM strain and orientation measurements via few-tilt precession, with potential for broad adoption across electron microscopy facilities.
Figure 1: MAPED experimental method and processing. (a) Diffraction patterns measured from five unique beam tilts. (b) Linear descan of the diffraction patterns as a function of probe position is corrected first. (c) After, real space field-of-view shifts are corrected across the five 4D-STEM experiments.
@@ 2.1 Data acquisition
An overview of our measurement scheme is shown in Fig. 1. First, we collected five 4D-STEM datasets, each at a different beam tilt; the beam was tilted using either the probe corrector (for aberration corrected instruments) or rotation center (in uncorrected instruments), shown in Fig. 1a. We measured and removed descan through computational post processing. This ensured that when summing up diffraction data, each pixel on every diffraction pattern corresponds to the same scattering angle of the electron beam. Tilting of the electron beam can also create shifts in the image plane as shown in Fig. 1c. These shifts are handled by image registration of virtual dark field images generated from each 4D-STEM dataset. Once alignment of the datasets converged, we averaged them and analyzed the average data in the same manner as conventional 4D-STEM data.
@@
The data in Fig. 2 were collected with a probe defined by a 10 $\mu$m custom bullseye aperture with a 2.2 mrad convergence semiangle  (Zeltmann et al., 2020). Five scans were acquired at indicated beam tilts of $(q_{x},q_{y})=(0,0)$, $(+9,0)$, $(-9,0)$, $(0,+9)$, and $(0,-9)$ mrad using the probe corrector (CESCOR, CEOS) beam tilt controls. After tilting the illumination beam, the diffraction pattern was manually shifted to the approximate center of the detector using post specimen deflectors. We scanned 256x256 probe positions with a dwell time of 500 µs. The control data were collected with 2500 µs dwell time with no added beam tilt. The electron fluence of both the tilt-averaged 4D-STEM dataset, and the control dataset was approximately 2.1 $\times$ 103e-12. Our sample was a MAG*I*CAL calibration sample available from Ted Pella, which consists of alternating silicon and silicon-germanium layers grown on a single crystal silicon substrate. The multilayers of Si and SiGe are grown with the [100] axis perpendicular to the interface; we chose this axis to be the x-axis for the strain maps. EELS data were collected using the Gatan K3 camera to calculate a thickness of t$/\lambda$ = 1.04 (Fig. S.1).
The data in Fig. 3 were acquired following a similar method as described for Fig. 2. Here a 5 $\mu$m custom aperture was used forming a probe with a 0.64 mrad convergence semiangle. Five scans were acquired with a dwell time of 100 µs at indicated beam tilts of $(q_{x},q_{y})=(0,0)$, $(+17,0)$, $(-17,0)$, $(0,+17)$, and $(0,-17)$ mrad using the probe corrector software. The control data was collected with 500 µs dwell time with no added beam tilt. The cumulative dose of both the control dataset and the tilt averaged dataset was approximately 2.8$\times$ 102e-12. The data were collected on the commercially available $\sim$30 nm thick evaporated aluminum diffraction standard sample from Ted Pella, suspended on a lacey carbon grid.
@@
The data in Fig. 5 were acquired on an uncorrected Thermo Fisher Scientific ThemIS operated in microprobe mode, where the illumination beam was tilted using the “rotation center” alignment feature of the microscope. By drawing a calibrated circle in the diffraction “View” mode and adjusting the rotation center until the unscattered beam intersected the circle, the corresponding rotation center X and Y values were recorded. These values varied approximately linearly with tilt between 9 and 17.5 mrad as shown in Fig. S.3. This calibration method does not account for a beam shift, but it is a quick and approximate approach. Data were collected using Gatan 4D-STEM software on a K2-IS camera operating at 400 frames per second. No diffraction shift correction was needed as the sensor was large enough to accommodate the pattern shift, making the entire workflow fast and straightforward. 4D-STEM data was also acquired on a Thermo Fisher Ceta camera using legacy TEM Imaging and Analysis (TIA) software. While effective, this workflow was more time-consuming, as the region of interest box had to be manually positioned and sized for each scan in the absence of scripting support. Diffraction patterns were acquired at 33 frames per second on the Ceta camera.
@@ 2.2 Data processing and analysis
As shown in Fig. 1b-c, the five-tilt series for both the strain and orientation mapping datasets were aligned in reciprocal space and then real space using the following procedure. First, we aligned the datasets in reciprocal space in order to ensure that every pixel on each diffraction pattern corresponds to the same scattering vector. After determining the center coordinate of each diffraction pattern (i.e, the coordinate of the unscattered beam) as a function of probe position, we aligned each pattern with subpixel accuracy to the center and resampled with bilinear interpolation. Each of the five datasets were aligned such that the unscattered beam was centered for all probe positions. These steps were performed using the calibration tools implemented in py4DSTEM (Savitzky et al., 2021). Finding the center beam can be challenging especially in a thick sample, where the brightest reflection is not necessarily the central spot, as exemplified by the K2 data shown in Fig. S.4. In this case, including a vacuum region in the scan or recording a subsequent image of the probe over vacuum with the same diffraction shift alignment can be helpful. If one does not adjust the unscattered beam with diffraction shift, the position of the center is a known value according to the calibration as shown in Fig. S.3 and this position value can be used as a first guess in the centering step.
Next, we calculated virtual dark-field images for each scan, and estimated the relative shifts in real space using cross correlation. For samples with fewer features, such as the SiGe standard, cross-correlation can be challenging, and it was often more effective to align tilted datasets to each other than to the centered one. Thus only one tilted dark-field image needed to be aligned to the center tilted image, which could be done automatically with the most similar tilted image or even manually if needed. The five datasets were then aligned in real space pixel-wise by rolling the datacube and then cropping to the region that contained data from all five scans. Fig. S.5 shows five of the tilts overlaid on top of each other.
We calculated strain and orientation maps using the workflows in py4DSTEM. For ACOM analysis of the MAPED dataset, the excitation error was modified using an estimate of 1, following previous work  (Diebold et al., 2025).
@@ 2.3 Simulations
We performed diffraction simulations in py4DSTEM using the dynamical diffraction module with the Weickenmeier-Kohl plus core and phonon absorption parameters at 300 kV with a maximum scattering angle of 4 Å-1. Instead of tilting the beam, the incident zone axis for the simulations was tilted accordingly. The average intensity across all tilts for the $\overline{1}\overline{1}0$ and $\overline{2}00$ are plotted in Fig. 4 for thickness ranging from 10 to 300 nm.
@@ 3.1 Strain mapping
Figure 2: (a) Single and (b) mean diffraction pattern from conventional experiment and corresponding (c–d) precession experiment. Strain maps from the (e) conventional and (f) precession experiment. Black arrows indicate location of alloy regions, where we expect large $\epsilon_{xx}$. Compared to precession, the conventional experiments show more error in the strain maps.
@@
Fig. 2 shows a comparison between strain mapping with the MAPED method and a conventional 4D-STEM experiment on the SiGe sample. In this sample, we expect lattice dilatation in the alloy layers compared to the pure silicon, as germanium has a larger lattice parameter than silicon (5.67 Å versus 5.43 Å) (Jain et al., 2013)). Because the y-direction is constrained due to the epitaxial growth, we expect the strain to be localized to the x-direction and the expansion of the lattice should create a tensile (positive) strain along that direction.
@@
These datasets were collected with a bullseye aperture to improve Bragg disk detection as described by  Zeltmann et al. (2020). Fig. 2a–b shows a single and mean diffraction pattern from a conventional experiment and Fig. 2c–d plots the corresponding MAPED patterns. The intensities of higher order Bragg reflections are clearly reduced in Fig. 2a–b, as a consequence of the curvature of the Ewald sphere. For the same dose, the data acquired with precession (Fig. 2c–d) shows that higher order Bragg reflections, which are excited upon tilting the beam. The improved signal to noise ratio for these higher order reflections enables us to more precisely locate them (via cross-correlation), and thus, improves the precision in our lattice constant measurements.
@@
Fig. 2e–f show the strain maps from the same field of view with and without MAPED, and the corresponding dark-field image is shown in Fig. S.6. A similar SiGe alloy sample has often been used as a benchmark for measuring strain with S/TEM techniques (Rouviere et al., 2013; Mahr et al., 2021; Munshi et al., 2022). The dataset collected with precession is much closer to the predicted strain values, as the $\epsilon_{xx}$ strain is uniform between and across different sets of alloy layers and shows tension due to the expanded lattice parameter. There are small fluctuations corresponding to the layers in $\epsilon_{yy}$, $\epsilon_{xy}$, and $\theta$, which could be real features of the material or artifacts. In comparison, the strain maps for the conventional scans show significant errors, including non-uniform strain between the sets of alloy layers, which is not expected for the identical SiGe layers. These errors are underscored by the histograms of strain from a control (non-alloy region) as shown in Fig. S.7. The distribution of strain intensities from an ideally control unstrained region is much higher without the MAPED method.
@@
There are multiple reasons why the strain mapping in the conventional experiment would be less reliable. First, a single tilt experiment is more susceptible to intensity variations in the disks leading to less accurate disk detection; comparing Fig. 2a&c, it is clear how precession leads to more uniform intensity across disks. Second, in the precession experiment, disks at higher scattering angles are more strongly excited and thus more clearly identified. The strain algorithm fits higher-order peaks in addition to first order reflections for more robust strain measurements, so the extra signal at high scattering angles can improve calculations. The strain measurements in both datasets could be further enhanced through advanced processing such as described by  Munshi et al. (2022). Ultimately, the combination of sequential precession and bullseye apertures leads to highly accurate strain measurements.

### v2

@@ Abstract
Precession of a converged beam during acquisition of a 4D-STEM dataset improves strain, orientation, and phase mapping accuracy by averaging over continuous angles of illumination. Precession experiments usually rely on integrated systems, where automatic alignments lead to fast, high-quality results. The dependence of these experiments on specific hardware and software is evident even when switching to non-integrated detectors on a precession tool, as experimental set-up becomes challenging and time-consuming. Here, we introduce multi-angle precession electron diffraction (MAPED): a method to perform electron diffraction by collecting sequential 4D-STEM scans at different incident beam tilts. The multiple diffraction datasets are averaged together post-acquisition, resulting in a single dataset that minimizes the impact of the curvature and orientation of the Ewald sphere relative to the crystal under study. Our results demonstrate that even four additional tilts improved measurement of material properties, namely strain and orientation, as compared to single-tilt 4D-STEM experiments. We show the versatility and flexibility of our MAPED approach with data collected on a number of microscopes with different hardware configurations and a variety of detectors.
@@ 1 Introduction
Figure 1: MAPED experimental method and processing. (a) Diffraction patterns measured from five unique beam tilts. (b) Descan of the diffraction patterns as a function of probe position is corrected first. (c) After, real space field-of-view shifts are corrected across the five 4D-STEM experiments.
@@
Modern STEM experiments often include 4D-STEM, where an electron beam is rastered across a specimen, and diffraction data are collected at each position in real space creating a 4-dimensional dataset (Ophus, 2019). The 4D-STEM data can be analyzed to map properties of materials that would be inaccessible or challenging to probe with conventional methods. For 4D-STEM strain experiments, data are collected with a probe formed by a small convergence angle, which reduces overlap of diffracted beams in reciprocal space. This approach can be extended to amorphous materials, which is not possible with conventional high resolution techniques (Gammer et al., 2018; Kennedy et al., 2025). Similarly, the orientation and phase of a sample can be studied using 4D-STEM, including automated crystal orientation mapping (ACOM) when coupled with computational diffraction pattern matching  (Rauch et al., 2010; Cautaerts et al., 2022; Ophus et al., 2022). Despite the immense benefits of strain mapping and ACOM techniques for studying the structure and properties of material systems, there can be practical challenges for thicker and more complex structures, due to multiple scattering of the electron beam leading to dynamical effects in the diffraction patterns (Müller et al., 2012).
@@
One of the most influential innovations for mitigating dynamical diffraction effects in S/TEM was introduced by Vincent and Midgley in 1994, who demonstrated precession electron diffraction with a convergent beam (Vincent and Midgley, 1994). This approach, which builds on earlier “hollow cone” illumination strategies  (Krakow and Howland, 1976; Kondo et al., 1984), uses rocking of the Ewald sphere to sample a wider range of diffraction space. The resulting diffraction patterns exhibit more uniform Bragg disk intensities, are less sensitive to sample mistilt, and can more easily excite higher-order reflections, making them more straightforward to index and more robust against dynamical scattering (Vincent and Midgley, 1994; Rauch et al., 2010; Rouviere et al., 2013). By averaging over multiple directions of incident electron beams, diffraction contrast can be suppressed, and it has even enabled the measurement of polarization through differential phase contrast (DPC) measurements (Kohno et al., 2022). Similar to this work, recent studies have implemented multiple tilts for analysis of electric polarization as an alternative to DPC imaging (Toyama et al., 2024; Flathmann et al., 2025).
@@
Precession electron diffraction can now be directly integrated into 4D-STEM workflows, where synchronized beam precession, raster scanning, and camera acquisition are handled by vendor-supplied systems. These integrated tools maintain precise alignment during acquisition, producing sharp diffraction patterns, which can be leveraged for high-quality analysis including strain mapping, orientation indexing, and digital dark field analysis (Corrêa et al., 2024; Diebold et al., 2025; MacLaren et al., 2025). These precession tools can be challenging to align, especially when they are used with hardware such as detectors and apertures that are not fully integrated in the system. Although there is an increasing emphasis and availability on precession tools, few microscopes still have integrated precession equipment, making the implementation of this technique limited.
@@
Here we present multi-angle precession electron diffraction (MAPED): a data-acquisition and post-processing strategy based on sequentially acquired 4D-STEM datasets at five distinct beam tilts to perform precession-like experiments. This sequential tilting approach allows us to replicate the averaging effect of hardware-integrated precession with the benefit of flexibility across different microscopes and detectors. We found that even four additional tilts ($\sim$0.5-1) dramatically improved the accuracy of the resulting strain maps. For challenging cases, such as phase discrimination in structurally similar crystals, we also simulated silicon samples along the [110] zone at various thicknesses to determine how well disk intensities matched those predicted by structure factor calculations. To test robustness, we examined data acquired on both corrected and uncorrected TEMs, using multiple camera systems, and assessed the effect of tilt count on indexing fidelity and strain reconstruction accuracy. We also provided open-source code such that this can be implemented at other microscopy centers. This work provides a generalizable and practical framework for enhancing 4D-STEM strain and orientation measurements, with potential for broad adoption across electron microscopy facilities.
@@ 2.1 Data acquisition
An overview of our measurement scheme is shown in Fig. 1. First, we collected five 4D-STEM datasets, each at a different beam tilt. We measured and removed descan through computational post processing. This ensured that when summing up diffraction data, each pixel on every diffraction pattern corresponds to the same scattering angle of the electron beam. Tilting of the electron beam can also create shifts in the image plane as shown in Fig. 1c. These shifts are handled by registration of virtual images generated from each 4D-STEM dataset. Once alignment of the datasets converged, we averaged them and analyzed the average data in the same manner as conventional 4D-STEM data.
@@
[…]
@@
The data in Fig. 5 were acquired on an uncorrected Thermo Fisher Scientific ThemIS at the Molecular Foundry operated in microprobe mode, where the illumination beam was tilted using the “rotation center” alignment feature of the microscope. By drawing a calibrated circle in the diffraction “View” mode and adjusting the rotation center until the unscattered beam intersected the circle, the corresponding rotation center X and Y values were recorded. These values varied approximately linearly with tilt between 9 and 17.5 mrad as shown in Fig. S.3. This calibration method does not account for a beam shift, but it is a quick and approximate approach. Data were collected using Gatan 4D-STEM software on a K2-IS camera operating at 400 frames per second. No diffraction shift correction was needed as the sensor was large enough to accommodate the pattern shift, making the entire workflow fast and straightforward. 4D-STEM data was also acquired on a Thermo Fisher Ceta camera using legacy TEM Imaging and Analysis (TIA) software. While effective, this workflow was more time-consuming, as the region of interest box had to be manually positioned and sized for each scan in the absence of scripting support. Diffraction patterns were acquired at 33 frames per second on the Ceta camera.
@@ 2.2 Data processing and analysis
As shown in Fig. 1b-c, the five-tilt series for both the strain and orientation mapping datasets were aligned in reciprocal space and then real space using the following procedure. First, we aligned the datasets in reciprocal space in order to ensure that every pixel on each diffraction pattern corresponds to the same scattering vector. After determining the center coordinate of each diffraction pattern (i.e, the coordinate of the unscattered beam) as a function of probe position, we aligned each pattern with subpixel accuracy to the center and resampled with bilinear interpolation. Each of the five datasets were aligned such that the unscattered beam was centered for all probe positions. These steps were performed using the calibration tools implemented in py4DSTEM (Savitzky et al., 2021), with more detail in Sec. S1. Finding the center beam can be challenging especially in a thick sample, where the brightest reflection is not necessarily the central spot, as exemplified by the K2-IS data shown in Fig. S.4. In this case, including a vacuum region in the scan or recording a subsequent image of the probe over vacuum with the same diffraction shift alignment can be helpful. If one does not adjust the unscattered beam with diffraction shift, the position of the center is a known value according to the calibration as shown in Fig. S.3 and this position value can be used as a first guess in the centering step.
Next, we calculated virtual dark-field images for each scan, and estimated the relative shifts in real space using cross correlation. For samples with fewer features, such as the SiGe standard, cross-correlation can be challenging, and it was often more effective to align tilted datasets to each other than to the centered one. Thus only one tilted dark-field image needed to be aligned to the center tilted image, which could be done automatically with the most similar tilted image or even manually if needed. The five datasets were then aligned in real space pixel-wise by rolling the datacube and then cropping to the region that contained data from all five scans. Fig. S.5 shows five of the tilts overlaid on top of each other. The datasets were only aligned with pixel-wise precision to reduce interpolation errors from multiple resampling of datasets. However, this reduces the real space resolution of the data, which could be limiting for some applications, in which case subpixel alignment would be appropriate.
We calculated strain and orientation maps using the workflows in py4DSTEM. The ACOM analysis of the MAPED dataset was performed using the sparse correlation matching method described in Ophus et al. (2022). The excitation error was modified to include a precession angle of 1, following previous work  (Diebold et al., 2025). After determining the best-fit orientation, we calculated the strain for each probe position using the following procedure. For each probe position, we simulated a reference diffraction pattern. Next, each simulated Bragg disk was paired with the closest experimentally measured Bragg disk, or removed if no disks were found within the radial threshold (set to 0.04 Å-1). We then used weighted least squares to calculate the best fit affine transformation matrix which deforms the reference pattern to match the experimental pattern. Finally, we plotted the infinitesimal strain from the transformation matrix for each probe position in the probe scanning coordinate system.
@@ 2.3 Simulations
We performed diffraction simulations in py4DSTEM using the dynamical diffraction module with the Weickenmeier-Kohl plus core and phonon absorption parameters at 300 kV with a maximum scattering angle of 4 Å-1. Instead of tilting the beam, the incident zone axis for the simulations was tilted accordingly. We simulate silicon with the [011] aligned with the electron beam, matching the experimental data in Fig. 2 for silicon. Fig. 4 shows the result of our calculations. We include one simulation at zero beam tilt, and N-1 equally spaced radially at 17 mrad. We performed simulations for sample thicknesses from 10 to 300 nm in 10 nm steps, using 1, 5, 9, 17, 33, 65, 129, 257, and 513 tilts.
@@ 3.1 Strain mapping
Figure 2: (a) Single and (b) mean diffraction pattern from conventional experiment and corresponding (c–d) MAPED experiment. Strain maps from the (e) conventional and (f) MAPED experiment. Black arrows indicate location of alloy regions, where we expect large $\epsilon_{xx}$. Compared to MAPED, the conventional experiments show more error in the strain maps.
@@
Fig. 2 shows a comparison between strain mapping with the MAPED method and a conventional 4D-STEM experiment on the SiGe sample. In this sample, we expect lattice dilatation in the alloy layers compared to the pure silicon, as germanium has a larger lattice parameter than silicon (5.67 Å versus 5.43 Å) (Jain et al., 2013). Because the y-direction is constrained due to the epitaxial growth, we expect the strain to be localized to the x-direction and the expansion of the lattice should create a tensile (positive) strain along that direction.
@@
These datasets were collected with a bullseye aperture to improve Bragg disk detection as described by  Zeltmann et al. (2020). Fig. 2a–b shows a single and mean diffraction pattern from a conventional experiment and Fig. 2c–d plots the corresponding MAPED patterns. The intensities of higher order Bragg reflections are clearly reduced in Fig. 2a–b, as a consequence of the curvature of the Ewald sphere. For the same dose, the data acquired with MAPED (Fig. 2c–d) shows that higher order Bragg reflections, which are excited upon tilting the beam.
@@
Fig. 2e–f show the strain maps from the same field of view with and without MAPED, and the corresponding dark-field image is shown in Fig. S.6. A similar SiGe alloy sample has often been used as a benchmark for measuring strain with S/TEM techniques (Rouviere et al., 2013; Mahr et al., 2021; Munshi et al., 2022). Here $\epsilon_{xx}$ corresponds with the x direction that is vertical to the page as indicated by the arrows to the left of the strain maps. Based on previous studies in the literature, we expect the strain to be approximately 1.2% (Rouviere et al., 2013; Munshi et al., 2022). The dataset collected with MAPED is much closer to the predicted strain values, as shown in Fig.2 and the line profiles in Fig. S.7. The $\epsilon_{xx}$ strain is close to uniform between and across different sets of alloy layers and shows tension due to the expanded lattice parameter. There are some additional features in the MAPED $\epsilon_{yy}$, $\epsilon_{xy}$, and $\theta$ strain maps, including shear strain across the field of view and small differences in strain between SiGe layers; some of these could be real features of our sample.
Nonetheless, in comparison, the strain maps for the conventional scans show significant errors, including non-uniform strain between the sets of alloy layers, which is not expected. These errors are underscored by the histograms of strain from a control (non-alloy region) as shown in Fig. S.8. The distribution of strain intensities from an ideally control unstrained region is much higher without the MAPED method.
@@
There are multiple reasons why the strain mapping in the conventional experiment would be less reliable. First, a single tilt experiment is more susceptible to intensity variations in the disks leading to less accurate disk detection; comparing Fig. 2a&c, it is clear how MAPED leads to more uniform intensity across disks. Second, in the MAPED experiment, disks at higher scattering angles are more strongly excited and thus more clearly identified. The strain algorithm fits higher-order peaks in addition to first order reflections for more robust strain measurements, so the extra signal at high scattering angles can improve calculations. The strain measurements in both datasets could be further enhanced through advanced processing such as described by  Munshi et al. (2022). Ultimately, the combination of sequential tilting and bullseye apertures leads to highly accurate strain measurements.